# Real-time Tracking in a Status Update System with an Imperfect Feedback Channel

### v1

@@ Abstract
We consider a status update system consisting of a finite-state Markov source, an energy-harvesting-enabled transmitter, and a sink. The forward and feedback channels between the transmitter and the sink are error-prone. We study the problem of minimizing the long-term time average of a (generic) distortion function subject to an energy causality constraint. Since the feedback channel is error-prone, the transmitter has only partial knowledge about the transmission results and, consequently, about the estimate of the source state at the sink. Therefore, we model the problem as a partially observable Markov decision process (POMDP), which is then cast as a belief-MDP problem. The infinite belief space makes solving the belief-MDP difficult. Thus, by exploiting a specific property of the belief evolution, we truncate the state space and formulate a finite-state MDP problem, which is then solved using the relative value iteration algorithm (RVIA). Furthermore, we propose a low-complexity transmission policy in which the belief-MDP problem is transformed into a sequence of per-slot optimization problems. Simulation results show the effectiveness of the proposed policies and their superiority compared to a baseline policy. Moreover, we numerically show that the proposed policies have switching-type structures.
Index Terms: Real-time tracking, status update, energy harvesting, Markov source, partially observable Markov decision process.
@@ I Introduction
Emerging applications in the Internet of Things (IoT) networks, such as digital health, intelligent transportation, and disaster monitoring, often rely on real-time tracking of a remotely monitored random process. In these systems, a sensor observes a physical process such as road traffic, patient heart rate, or temperature and transmits real-time status updates to an intended destination, such as a remote controller or a decision-making entity [1].
@@
A commonly used metric to quantify the information freshness in status update systems is the Age of Information (AoI) [2, 3]. AoI is defined as the difference between the current time and the generation time of the last received status update at the destination [2, 3]. Extensive research has been conducted on the AoI analysis and optimization in different areas, e.g., queuing systems [4, 5], scheduling [6, 7], and sampling problems [8, 9]. However, AoI solely focuses on the time elapsed since the last received update’s generation, neglecting the updates’ semantic content and the mismatch between the current value of the process of interest and its estimate at the destination [10, 11]. This motivated us to address the problem of optimizing (generic) distortion measures for real-time tracking in status update systems [11].
@@
As the IoT and wireless sensor networks expand, traditional battery-powered sensors struggle to meet the growing demand for sustainability and low maintenance. Energy harvesting (EH) technology offers a promising alternative, allowing sensors to draw power from their environment, eliminating battery replacements, and significantly extending lifespans [12]. For instance, EH sensors can utilize sunlight, vibrations, thermal energy, or even ambient radio waves to harvest energy [12].
@@
In this work, we consider a status update system where an EH-enabled transmitter monitors a source and sends status updates to a sink over an unreliable wireless channel (see Fig. 1). The source is modeled as a finite-state symmetric Markov chain. The transmitter is equipped with a battery with limited capacity that is charged by harvesting energy from the environment. We assume a slotted communication system in which, after each successful reception, the sink sends an acknowledgment signal to the transmitter over a noisy feedback channel. Our considered system may model a part of intelligent transportation systems in which an energy-harvesting roadside unit monitors the status of the vehicle flows and sends updates to a traffic management center over a wireless communication channel [13, 14].
@@
We address a real-time tracking problem, where the main goal is to minimize the expected time average of a (generic) distortion function subject to an energy causality (availability) constraint. The distortion function measures the discrepancy between the source state and its estimated value at the sink (e.g., the absolute value of the difference between the source state and its estimated value at the sink or an indicator function representing the mismatch between the source and the sink). The solution to the problem determines the optimal transmission times. Since the feedback channel is error-prone, some of the acknowledgments may not be detected at the transmitter. Thus, the sink’s estimate of the source is only partially observable at the transmitter side. Thus, we model our problem as a partially observable Markov decision process (POMDP), and subsequently, we cast it as a belief-MDP problem. Due to infinite belief space, solving the original belief-MDP problem is difficult. Thus, we truncate the state space by exploiting a specific property of the belief evolution. By truncation, we formulate a finite state belief-MDP problem, which is then solved using the relative value iteration algorithm (RVIA). Since the complexity of the RVIA-based policy increases with the number of source states and the battery’s capacity, we propose a sub-optimal low-complexity transmission policy. The main idea of this policy is that instead of solving the original belief-MDP problem, the problem is transformed into a sequence of per-slot optimization problems. In the numerical results, we study the effectiveness of the proposed policies and compare the proposed policies with a baseline policy. Moreover, we study the structure of proposed policies, which reveals their switching-type structure.
@@
  • We study a real-time tracking problem in a status update system with an imperfect feedback channel. We investigate the problem of minimizing the average distortion function under an energy causality constraint.
  • We model the main problem as a POMDP and propose two transmission policies: 1) a deterministic transmission policy by truncating the belief space and then applying the RVIA, and 2) a sub-optimal low-complexity policy by transforming the belief-MDP problem into a sequence of per-slot problems.
  • We numerically evaluate the proposed policies’ effectiveness under different settings. We also investigate the structure of proposed policies and show their switching-type structure.
@@ I-A Related Work
Real-time tracking and remote state estimation have recently witnessed great interest in numerous studies, e.g., [15, 16, 17, 18, 11, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. The authors of [15] considered a hybrid automatic repeat request (HARQ) based real-time remote estimation framework for a status update system. They proposed an optimal scheduling policy to minimize the long-term mean squared error (MSE). The work [16] considered a two-way delay status update system where the controller is at the sink and the source is modeled as the Wiener process. They developed a jointly optimal sensor/controller waiting policy to minimize estimation error from the sensor perspective and AoI from the controller perspective. In [17], the authors proposed an optimal quantization strategy for real-time monitoring of an Ornstein-Uhlenbeck (OU) process in a status update system. They minimized the time-average MSE subject to an AoI outage constraint. The authors of [18] studied a status update system with temporally and spatially correlated source states. They used a time-varying Gauss-Markov random field to model the source state and derived the closed-form expression of the average estimation error. The work [11] studied a status update system where the source is modeled as a Wiener process and proposed an optimal online sampling strategy to minimize the MSE. They studied the connection between the policy that minimizes the AoI and the policy that minimizes the remote estimation error. In [19], the authors generalized the work in [11] by considering a source modeled as OU process. They showed that the optimal sampling policy has a switching-type structure. The authors of [20] investigated the real-time remote estimation in a status update system with a HARQ protocol and different retransmission policies. They proposed an optimal transmission control policy to minimize the average MSE and a variation of MSE. The work [21] addressed the problem of remote estimation in a multi-source status update system. Using the Whittle index theory, they proposed a heuristic channel allocation policy. This policy aims to find an optimal tradeoff between the total discounted estimation error and the energy consumption over the infinite horizon. In [22], the authors studied the real-time reconstruction problem of a continuous process at a remote destination. They derived the closed-form expression of the average distortion between the original and the estimated process. Moreover, they found the optimal sampling rate by minimizing the derived average distortion. The authors of [23] studied the problem of minimizing the weighted sum of the expected time-average estimation errors of sources in a status update system. They proposed a sampling and transmission scheduling policy to solve the problem using the Whittle index theory. The work [24] studied the problem of remote estimation of a source modeled as the Wiener process over a channel with unknown delay statistics. They proposed an online sampling policy to minimize the expected time average MSE under a sampling frequency constraint. In [25], the authors derived a closed-form expression for the average MSE under two different scheduling policies in a multi-source status update system. The authors of [26] considered a status update system consisting of a vehicle as the source and a roadside unit as the controller. They proposed a scheduling policy to minimize the average weighted estimation error under a resource constraint. The work [27] considered a multi-source status update system and addressed the problem of joint estimation error and transmission cost minimization. They considered two threshold-based policies and evaluated them numerically.
@@
The most related works to this paper are [28, 29, 30]. Similar to our work, the authors of [28] studied a status update system in which the source dynamics is molded via a Markov chain, and the transmitter harvests energy for operation. However, they considered a perfect communication channel and solved the real-time estimation problem for a finite time horizon. Considering an unreliable channel (both forward and feedback channels) makes the system more complicated as the status updates are not delivered for all transmission attempts, and the transmitter does not always know whether the sent status updates are successfully received. The work [29] considered a similar real-time tracking system yet with the following fundamental differences. The authors in [29] considered perfect feedback, whereas we consider imperfect feedback which makes the sink only partially observable to the transmitter. In addition, they considered a binary Markov source, while we consider a multi-state source. In [30], the authors considered a status update system with noisy channels in the forward and feedback channels. Unlike [30], we consider an energy harvesting powered transmitter which makes a complicated system as the transmitter does not know the availability of the energy at the subsequent slots. Moreover, the work [30] studied the AoI in the system, which does not consider the sample’s value, whereas we study a real-time tracking problem.
I-B Organization
The rest of this paper is organized as follows. The system model and problem formulation are presented in Section II. The proposed policies are presented in Section III. Numerical results are presented in Section IV. Finally, concluding remarks are made in Section V.
@@ II-A System Model
The source is modeled as a finite-state symmetric discrete-time Markov chain [^1] This is a commonly used source model (see e.g., [28, 31, 32]). and is fully observable by the transmitter. Let ${X(t)\in\{1,\ldots,N\}}$ denote the source state at slot $t$, where $N$ is the number of source states. Each slot, the source either remains in the same state with a probability $p$ or transits [^2] We assume that the source and the system clocks are synchronized. to any other state with a probability $q=(1-p)/(N-1)$, as shown in Fig. 1. Let $P$ denote the state transition matrix of the source, given as
Note that $P$ is a doubly stochastic matrix, i.e., a square matrix with non-negative entries between 00 and $1$, where each row and column sums up to $1$.
@@
The transmitter sends the status update over an error-prone wireless channel. We denote the probability of successful reception by $p_{s}$. After each successful reception, the sink sends an acknowledgment (ACK) to the transmitter. We assume that the feedback channel is delay-free but error-prone, i.e., ACK is either received instantaneously or lost. Let $p_{f}$ denote the probability of a successful ACK reception over the feedback channel. Let $d(t)\in\{0,1\}$ be the binary indicator of the reception of ACK at slot $t$; $d(t)=1$ indicates that the transmitter receives an ACK signal after a transmission attempt, and $d(t)=0$ indicates that the transmitter does not receive an ACK due to either an error over the forward channel (i.e., the status update is not received by the sink) or due to an error over the feedback channel (i.e., status update is correctly received at the sink but the ACK is lost). We also define $d(t)=0$ for the slots where the transmitter sends no status update.
@@
See appendix A
At each slot $t$, the transmitter decides whether to stay idle or to send a new status update based on the following available information: 1) the current source state $X(t)$, 2) the battery’s energy level $b(t)$, 3) the previous source state $X(t-1)$, and 4) the ACK reception indicator at the previous slot ${d(t-1)}$.
@@ II-B Problem Formulation
Our main goal is to minimize the average estimation error, i.e., the discrepancy between the source state $X(t)$ and its estimated value at the sink $\hat{X}(t)$, by optimizing transmission decisions subject to the energy limitation imposed by the energy harvesting circuit.
Let $g:\big{(}X,\hat{X}\big{)}\rightarrow\mathbb{R}^{+}$ denote a (generic) bounded distortion function, i.e., $|g(.)|<\infty$, meant to quantify the estimation error. For instance, the distortion function can represent the absolute value of the difference between $X(t)$ and $\hat{X}(t)$, i.e., ${g\big{(}X,\hat{X}\big{)}=|X-\hat{X}|}$. Then, the considered problem can be formulated as the following stochastic control problem:
@@
with variables $a(t)$ for all $t\in\{1,2,\dots\}$. Inequality (4b) represents the energy causality constraint, where the transmitter can send a status update if the battery level is larger than zero. The expression (4c) indicates the binary nature of the decision variable. In problem (4), $\mathbb{E}\{\cdot\}$ is the expectation with respect to the randomness of the system (i.e., randomness in the source dynamic, the energy arrival, and the forward and feedback communication channels) and the decision variable $a(t)$.
@@ III-A POMDP Formulation
States: Let ${s(t)=\big{(}X(t),b(t),\hat{X}(t),d(t-1),X(t-1)\big{)}\in\mathcal{S}}$ denote the state of the system at slot $t$, where $\mathcal{S}$ is the state space. Note that, due to the imperfect feedback channel, the estimated value of the source state at the sink $\hat{X}(t)$ is only partially observable at the transmitter side.
Actions: Let $a(t)\in\mathcal{A}_{s}$ denote the action at slot $t$, where $\mathcal{A}_{s}$ denotes the space of the feasible actions at state $s$, defined as
$\mathcal{A}_{s}=\big{\{}a\in\{0,1\}\mid b-a\geq 0\big{\}}.$ (5)
Observations: Let ${o(t)=\big{(}X(t),b(t),d(t-1),X(t-1)\big{)}}$ denote the observation at slot $t$, i.e., the observable part of the state. In general, the observation function [33, Chapter 7] is defined as the probability distribution function of the current observation given the current state and previous action, i.e., ${\mathrm{Pr}(o(t)\mid s(t),a(t-1))}$. Since the observation is always part of the state, the observation function can be expressed as
$\mathrm{Pr}\big{(}o(t)|s(t),a(t-1)\big{)}=\mathbbm{1}_{\{o(t)=\left(X(t),b(t),d(t-1),X(t-1)\right)\}},$ (6)
[…]
Cost: Let ${C\big{(}s(t)\big{)}=g\big{(}X(t),\hat{X}(t)\big{)}}$ denote the cost at slot $t$, where $g(\cdot)$ is the distortion function.
State transition probabilities: Let ${\mathrm{Pr}(s^{\prime}\mid s,a)}$ denote the probability of moving from the current state ${s}$ to the next state ${s^{\prime}}$ under action ${a}$. For the sake of brevity, hereinafter, we denote $\bar{\mu}=1-\mu$, $\bar{p}_{f}=1-p_{f}$, $\bar{p}_{s}=1-p_{s}$, and ${\tilde{b}=\min(b+1,B)}$. We further use $X_{-1}$ to denote the source state at the previous slot, $d_{-1}$ to denote the indicator of reception of ACK at the previous slot, ${s=(X,b,\hat{X},d_{-1},X_{-1})}$, and ${s^{\prime}=(X^{\prime},b^{\prime},\hat{X}^{\prime},d_{-1}^{\prime},X_{-1}^{\prime})}$. With these notations, the state transition probabilities can be expressed as
$\mathrm{Pr}(s^{\prime}\mid s,a)=$
$\begin{cases}p\bar{\mu}&a=0,s^{\prime}=(X,b,\hat{X},0,X)\\ p\mu&a=0,s^{\prime}=(X,\tilde{b},\hat{X},0,X)\\ q\bar{\mu}&a=0,s^{\prime}=(X^{\prime},b,\hat{X},0,X)\\ q\mu&a=0,s^{\prime}=(X^{\prime},\tilde{b},\hat{X},0,X)\\ p\bar{\mu}p_{s}p_{f}&a=1,s^{\prime}=(X,b-1,X,1,X)\\ p\mu p_{s}p_{f}&a=1,s^{\prime}=(X,b,X,1,X)\\ q\bar{\mu}p_{s}p_{f}&a=1,s^{\prime}=(X^{\prime},b-1,X,1,X)\\ q\mu p_{s}p_{f}&a=1,s^{\prime}=(X^{\prime},b,X,1,X)\\ p\bar{\mu}p_{s}\bar{p}_{f}&a=1,s^{\prime}=(X,b-1,X,0,X)\\ p\mu p_{s}\bar{p}_{f}&a=1,s^{\prime}=(X,b,X,0,X)\\ q\bar{\mu}p_{s}\bar{p}_{f}&a=1,s^{\prime}=(X^{\prime},b-1,X,0,X)\\ q\mu p_{s}\bar{p}_{f}&a=1,s^{\prime}=(X^{\prime},b,X,0,X)\\ p\bar{\mu}\bar{p}_{s}&a=1,s^{\prime}=(X,b-1,\hat{X},0,X)\\ p\mu\bar{p}_{s}&a=1,s^{\prime}=(X,b,\hat{X},0,X)\\ q\bar{\mu}\bar{p}_{s}&a=1,s^{\prime}=(X^{\prime},b-1,\hat{X},0,X)\\ q\mu\bar{p}_{s}&a=1,s^{\prime}=(X^{\prime},b,\hat{X},0,X)\\ 0&\text{otherwise},\end{cases}$ (7)

### v2

@@ Abstract
We consider a status update system consisting of a finite-state Markov source, an energy-harvesting-enabled transmitter, and a sink. The forward and feedback channels are error-prone. We study the problem of minimizing the long-term time average of a (generic) distortion function subject to an energy limitation. Since the feedback channel is error-prone, the transmitter has only partial knowledge about the transmission results and, consequently, about the estimate of the source state at the sink. Therefore, we model the problem as a partially observable Markov decision process (POMDP), which is then cast as a belief-MDP problem. The infinite belief-state space makes solving the belief-MDP difficult. Thus, by exploiting a specific property of the belief evolution, we truncate the state space and formulate a finite-state MDP problem, which is then solved using the relative value iteration algorithm (RVIA). Furthermore, we propose an energy-agnostic low-complexity policy in which the belief-MDP problem is transformed into a sequence of per-slot optimization problems. Then, the energy-agnostic low-complexity policy is extended to an energy-aware low-complexity policy by adding a regularization term to the objective function of the per-slot problems. Simulation results show the structure and effectiveness of the proposed policies and their superiority compared to baseline policies.
Index Terms: Real-time tracking, status updating, energy harvesting, partially observable Markov decision process.
@@ I Introduction
Emerging applications in the Internet of Things (IoT) networks, such as digital health, intelligent transportation, and disaster monitoring, often rely on real-time tracking of a remotely monitored random process. In these systems, a sensor observes a physical process such as road traffic, patient heart rate, or temperature, and transmits real-time status updates to an intended destination, such as a remote controller or a decision-making entity [2].
@@
A commonly used metric to quantify the information freshness in status update systems is the age of information (AoI) [3, 4]. AoI is defined as the difference between the current time and the generation time of the last received status update at the destination [3, 4]. Extensive research has been conducted on the AoI analysis and optimization in different areas, e.g., queuing systems [5], scheduling [6], and sampling problems [7, 8]. However, AoI solely focuses on the time elapsed since the last received update’s generation, neglecting the updates’ content and the mismatch between the current value of the process of interest and its estimate at the destination [9, 10, 11, 12]. This motivated us to address the problem of optimizing (generic) distortion measures for real-time tracking in status update systems. The distortion function is meant to capture the discrepancy between the monitored process and its estimate at the sink relative to the application’s goal. Notably, the distortion here can be tailored to error indicator [13], absolute error[14], or squared error [15].
To make optimal decisions in a status update system, the controller must know whether the sent packet was successfully delivered. To this end, the receiver is required to send acknowledgment feedback to the controller through a wireless channel upon a status update delivery. Ensuring an error-free feedback channel is challenging due to noise, interference, and fading. To address this challenge, we study a status update system with an imperfect feedback channel.
@@
Furthermore, as the IoT and wireless sensor networks expand, traditional battery-powered sensors struggle to meet the growing demand for sustainability and low maintenance costs. Energy harvesting (EH) technology offers a promising alternative, allowing sensors to draw power from their environment, eliminating battery replacements, and significantly extending their lifespans [16].
@@
In this work, we consider a status update system where an EH-enabled transmitter monitors a source and sends status updates to a sink over an unreliable wireless channel (see Fig. 1). The source is modeled as a finite-state Markov chain. The transmitter is equipped with a battery with limited capacity that is charged by harvesting energy from the environment. We assume a slotted communication system in which, after each successful reception, the sink sends an acknowledgment signal to the transmitter over an imperfect feedback channel.
The considered system model may model status updating in an intelligent transportation system that uses EH-based roadside units [17, 18, 19]. These units harvest renewable energy sources, such as solar or wind energy, to sustain their operations. They monitor critical parameters, such as traffic flow and vehicular density, and transmit the collected data to a traffic management center via a wireless communication channel. Given the nature of such applications, the destination is interested in the real-time tracking of the monitored parameters [17, 18, 19].
@@
We address a real-time tracking problem, where the goal is to minimize the expected time average of a (generic) distortion function subject to an energy limitation constraint imposed by the energy harvesting circuit. The solution to the problem determines the optimal transmission times. Since the feedback channel is error-prone, some of the acknowledgments may not be detected at the transmitter. Thus, the sink’s estimate of the source is only partially observable at the transmitter side. Thus, we model our problem as a partially observable Markov decision process (POMDP), and subsequently, we cast it as a belief-MDP problem. Due to infinite belief space, solving the original belief-MDP problem is difficult. Thus, we truncate the state space by exploiting a specific property of the belief evolution. By truncation, we formulate a finite state belief-MDP problem, which is then solved using the relative value iteration algorithm (RVIA). Since the complexity of the POMDP-based policy increases with the number of source states and the battery’s capacity, we propose two sub-optimal low-complexity (LC) transmission policies, namely: 1) energy-agnostic LC policy and 2) energy-aware LC policy. The main idea behind the energy-agnostic LC policy is that instead of solving the original belief-MDP problem, the problem is transformed into a sequence of per-slot optimization problems. When the transmitter has both options of transmitting and staying idle, the energy-agnostic LC policy makes decisions independently of the energy arrival process, while the energy arrival process affects the long-term energy availability for the transmitter’s operations. Thus, by adding a regularization term to the objective function of the per-slot problems, we propose the energy-aware LC policy. The regularization term is based on the energy arrival process and promotes less transmission to save energy for future use. Specifically, the regularization term decreases the transmission rate when the energy arrival rate decreases, and its effect decreases when the energy arrival rate increases. In the numerical results, we study the effectiveness of the proposed policies and compare the proposed policies with a baseline policy. Moreover, we study the structure of proposed policies, which reveals their switching-type structure.
@@
  • We address a real-time tracking problem in a status update system with an imperfect feedback channel. We investigate the problem of minimizing an average distortion function under an energy constraint imposed by the energy harvesting circuit.
  • We model the main problem as a POMDP and propose three transmission policies: 1) a deterministic transmission policy by truncating the belief space and then applying the RVIA, 2) a sub-optimal energy-agnostic LC policy by transforming the belief-MDP problem into a sequence of per-slot problems, and 3) a sub-optimal energy-aware LC policy that takes the energy arrival rate into consideration.
  • We numerically evaluate the proposed policies’ effectiveness under different settings. We also investigate the structure of the proposed policies and show their switching-type structure.
@@ I-A Related Work
Real-time tracking and remote state estimation have recently witnessed great interest in numerous studies, e.g., [15, 20, 21, 22, 12, 23, 24, 14, 25, 26, 27, 28, 29, 30, 31, 32]. The work [14] addressed the problem of remote estimation in a multi-source status update system. Using the Whittle index theory, they proposed a heuristic channel allocation policy. This policy aims to find an optimal tradeoff between the total discounted estimation error and the energy consumption over the infinite horizon. The authors of [15] considered a hybrid automatic repeat request (HARQ) based real-time remote estimation framework for a status update system. They proposed an optimal scheduling policy to minimize the long-term mean squared error (MSE). The work [20] considered a two-way delay status update system where the controller is at the sink and the source is modeled as the Wiener process. They developed a jointly optimal sensor/controller waiting policy to minimize estimation error from the sensor perspective and AoI from the controller perspective. In [21], the authors proposed an optimal quantization strategy for real-time monitoring of an Ornstein-Uhlenbeck (OU) process in a status update system. They minimized the time-average MSE subject to an AoI outage constraint. The authors of [22] studied a status update system with temporally and spatially correlated source states. They used a time-varying Gauss-Markov random field to model the source state and derived the closed-form expression of the average estimation error. The work [12] studied a status update system where the source is modeled as a Wiener process and proposed an optimal online sampling strategy to minimize the MSE. They studied the connection between the policy that minimizes the AoI and the policy that minimizes the remote estimation error. In [23], the authors generalized the work in [12] by considering a source modeled as OU process. They showed that the optimal sampling policy has a switching-type structure. The authors of [24] investigated the real-time remote estimation in a status update system with a HARQ protocol and different retransmission policies. They proposed an optimal transmission control policy to minimize the average MSE and a variation of MSE. In [25], the authors studied the real-time reconstruction problem of a continuous process at a remote destination. They derived the closed-form expression of the average distortion between the original and the estimated process. Moreover, they found the optimal sampling rate by minimizing the derived average distortion. The authors of [26] studied the problem of minimizing the weighted sum of the expected time-average estimation errors of sources in a status update system. They proposed a sampling and transmission scheduling policy to solve the problem using the Whittle index theory. The work [27] studied the problem of remote estimation of a source modeled as the Wiener process over a channel with unknown delay statistics. They proposed an online sampling policy to minimize the expected time average MSE under a sampling frequency constraint. In [28], the authors derived a closed-form expression for the average MSE under two different scheduling policies in a multi-source status update system. The authors of [29] considered a status update system consisting of a vehicle as the source and a roadside unit as the controller. They proposed a scheduling policy to minimize the average weighted estimation error under a resource constraint. The work [30] considered a multi-source status update system and addressed the problem of joint estimation error and transmission cost minimization. They considered two threshold-based policies and evaluated them numerically. The work [31] considered an energy harvesting status update system where the source is only partially observable. They proposed several scheduling policies to minimize average distortion and age of incorrect information.
@@
The most related works to this paper are [32, 33]. Similar to our work, the authors of [32] studied a status update system in which the source dynamic is molded via a Markov chain, and the transmitter harvests energy for operation. They considered a perfect communication channel and solved the real-time estimation problem for a finite time horizon. Different from [32], we consider imperfect forward and feedback channels and a long-term average-time problem, which makes our problem more complicated such that the solution of [32] is not applicable to solve it. Notably, due to imperfect channels, status updates are not delivered for all transmission attempts, and the transmitter does not always know whether the sent status updates are successfully received. In [33, 34], the authors considered a status update system with an imperfect feedback channel. However, in [33, 34], the authors studied a different problem; we solve a control optimization problem to minimize the average distortion function whereas they characterized a closed-form expression of the average AoI under a fixed policy. Moreover, we consider an energy-harvesting powered transmitter, which makes the system more complicated as the transmitter does not know the availability of the energy at the subsequent slots.
@@ II-A System Model
The source is modeled as a finite-state symmetric discrete-time Markov chain [^1] This is a commonly used source model (see e.g., [32, 35, 36]). and is fully observable by the transmitter. Let ${X(t)\in\{1,\ldots,N\}}$ denote the source state at slot $t$, where $N$ is the number of source states. Each slot, the source either remains in the same state with a probability $p$ or transits [^2] We assume that the source and the system clocks are synchronized. to any other state with a probability $q=(1-p)/(N-1)$, as shown in Fig. 1.
@@
The transmitter sends the status update over an error-prone wireless channel. We denote the probability of successful reception by $p_{s}$. After each successful reception, the sink sends an acknowledgment (ACK) to the transmitter. We assume that the feedback channel is delay-free but error-prone, i.e., ACK is either received instantaneously or lost. Let $p_{f}$ denote the probability of a successful ACK reception over the feedback channel. Let $f(t)\in\{0,1\}$ be the binary indicator of the reception of ACK at slot $t$; $f(t)=1$ indicates that the transmitter receives an ACK signal after a transmission attempt, and $f(t)=0$ indicates that the transmitter does not receive an ACK due to either an error over the forward channel (i.e., the status update is not received by the sink) or due to an error over the feedback channel (i.e., status update is correctly received at the sink but the ACK is lost). We also define $f(t)=0$ for the slots where the transmitter sends no status update.
@@
See appendix A. ∎
At each slot $t$, the transmitter decides whether to stay idle or to send a new status update based on the following available information: 1) the current source state $X(t)$, 2) the battery’s energy level $b(t)$, 3) the previous source state $X(t-1)$, and 4) the ACK reception indicator at the previous slot ${f(t-1)}$.
@@ II-B Problem Formulation
Our main goal is to minimize the average estimation error, i.e., the discrepancy between the source state $X(t)$ and its estimated value at the sink $\hat{X}(t)$, by optimizing transmission decisions subject to an energy limitation imposed by the energy harvesting circuit.
[…]
@@
with variables $a(t)$ for all $t\in\{1,2,\dots\}$. Inequality (3b) represents the energy limitation constraint, where the transmitter can send a status update if the battery level is larger than zero. The expression (3c) indicates the binary nature of the decision variable. In problem (3), $\mathbb{E}\{\cdot\}$ is the expectation with respect to the randomness of the system (i.e., randomness in the source dynamic, the energy arrival, and the forward and feedback communication channels) and the decision variable $a(t)$.
@@ III-A POMDP Formulation
States: Let ${s(t)=\big{(}X(t),b(t),\hat{X}(t),f(t-1),X(t-1)\big{)}\in\mathcal{S}}$ denote the state of the system at slot $t$, where $\mathcal{S}$ is the state space. Note that, due to the imperfect feedback channel, the estimated value of the source state at the sink $\hat{X}(t)$ is only partially observable at the transmitter side.
Actions: Let $a(t)\in\mathcal{A}_{s(t)}$ denote the action at slot $t$, where $\mathcal{A}_{s}$ denotes the space of the feasible actions at state $s$, defined as $\mathcal{A}_{s}=\big{\{}a\in\{0,1\}\mid b-a\geq 0\big{\}}$.
Observations: Let ${o(t)=\big{(}X(t),b(t),f(t-1),X(t-1)\big{)}}$ denote the observation at slot $t$, i.e., the observable part of the state. In general, the observation function [37, Chapter 7] is defined as the probability distribution function of the current observation given the current state and previous action, i.e., ${\mathrm{Pr}(o(t)\mid s(t),a(t-1))}$. Since the observation is always part of the state, the observation function can be expressed as
$\mathrm{Pr}\big{(}o(t)|s(t),a(t-1)\big{)}=\mathbbm{1}_{\{o(t)=\left(X(t),b(t),f(t-1),X(t-1)\right)\}},$ (4)
Cost: Let ${C\big{(}s(t)\big{)}=d\big{(}X(t),\hat{X}(t)\big{)}}$ denote the cost at slot $t$, where $d(\cdot)$ is the distortion function.
State transition probabilities: Let $\mathcal{P}(s^{\prime}\mid s,a)$ denote the state transition probability defined as the probability of moving from the current state ${s}$ to the next state ${s^{\prime}}$ under action ${a}$. For the sake of brevity, hereinafter, we denote $\bar{\mu}=1-\mu$, $\bar{p}_{f}=1-p_{f}$, $\bar{p}_{s}=1-p_{s}$, and ${\tilde{b}=\min(b+1,B)}$. We further use $X_{-1}$ to denote the source state at the previous slot, $f_{-1}$ to denote the indicator of reception of ACK at the previous slot, ${s=(X,b,\hat{X},f_{-1},X_{-1})}$, and ${s^{\prime}=(X^{\prime},b^{\prime},\hat{X}^{\prime},f_{-1}^{\prime},X_{-1}^{\prime})}$. With these notations, the state transition probabilities can be expressed as
$\mathcal{P}(s^{\prime}\mid s,a)=\begin{cases}p\bar{\mu}&a=0,s^{\prime}=(X,b,\hat{X},0,X)\\ p\mu&a=0,s^{\prime}=(X,\tilde{b},\hat{X},0,X)\\ q\bar{\mu}&a=0,s^{\prime}=(X^{\prime},b,\hat{X},0,X)\\ q\mu&a=0,s^{\prime}=(X^{\prime},\tilde{b},\hat{X},0,X)\\ p\bar{\mu}p_{s}p_{f}&a=1,s^{\prime}=(X,b-1,X,1,X)\\ p\mu p_{s}p_{f}&a=1,s^{\prime}=(X,b,X,1,X)\\ q\bar{\mu}p_{s}p_{f}&a=1,s^{\prime}=(X^{\prime},b-1,X,1,X)\\ q\mu p_{s}p_{f}&a=1,s^{\prime}=(X^{\prime},b,X,1,X)\\ p\bar{\mu}p_{s}\bar{p}_{f}&a=1,s^{\prime}=(X,b-1,X,0,X)\\ p\mu p_{s}\bar{p}_{f}&a=1,s^{\prime}=(X,b,X,0,X)\\ q\bar{\mu}p_{s}\bar{p}_{f}&a=1,s^{\prime}=(X^{\prime},b-1,X,0,X)\\ q\mu p_{s}\bar{p}_{f}&a=1,s^{\prime}=(X^{\prime},b,X,0,X)\\ p\bar{\mu}\bar{p}_{s}&a=1,s^{\prime}=(X,b-1,\hat{X},0,X)\\ p\mu\bar{p}_{s}&a=1,s^{\prime}=(X,b,\hat{X},0,X)\\ q\bar{\mu}\bar{p}_{s}&a=1,s^{\prime}=(X^{\prime},b-1,\hat{X},0,X)\\ q\mu\bar{p}_{s}&a=1,s^{\prime}=(X^{\prime},b,\hat{X},0,X)\\ 0&\text{otherwise},\end{cases}$ (5)